### v5

# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

Abstract
This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for target search in unknown environments. We consider the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings. These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks. The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier. An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target. In order to achieve this, we propose a utility function that accounts for traditional metrics like information gain and path cost and also for the additional contextual information from the scan classifier. LIVES is baselined against several existing exploration methods in simulation to verify its performance. Finally, it is validated in real-world experiments searching for single and multiple targets with a Spot robot in two unseen environments. Videos of experimental validation, implementation details and open source code can be found on our project website at https://sites.google.com/view/lives-2024/home.
I INTRODUCTION
Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [35], target search [1, 12], disaster response [15, 40, 4], agricultural monitoring [13], reconnaissance and surveillance [16, 20, 23]. Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. This requires planners equipped to generate high quality plans under uncertainty. Conventional methods for LiDAR and multi sensor exploration effectively map environments with focus on completeness, rather than being tailored time critical missions, which we investigate in this paper.
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. Use of scans is often limited to obstacle avoidance, localization and mapping. In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing. While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17]. Real time projection-based methods instead aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26]. There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43]. There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation. This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
Figure 1: Overview of a planning step of LIVES. In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot. They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c). Separately, LiDAR scans and pose data are used to maintain the global search maps (d). Global search maps are used to generate frontier points (e), which are used for finding candidate viewpoints. Finally, each frontier point is sampled at four viewpoints (f) and scored III-E2. In this toy example the left viewpoint in (f) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed. Finally, the selected viewpoint is sent back to the robot for execution.
The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target. In particular, this method deviates from traditional exploration methods, which focus on completeness. Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings. We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information. We accomplish this using a map based classifier to train a model to recognize features inherent to many structured, indoor environments. The perception model feeds the planner those points unlikely to be permanent, non-map points, for use during viewpoint planning. These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment. Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points. This information obviates the need to visually inspect parts of the environment that are likely to be permanent. An overview of the proposed method can be found in Fig. 1.
The contributions of this work can be summarized as follows:
  • A learning strategy for online scan segmentation as map and non-map in unseen, real-world environments
  • A formulation that incorporates classified scan pixel information into a planner for visual search
  • Demonstrate reduced mission completion time in the visual search task over existing planners
  • Deploy and verify the proposed method on a real robotic system in multiple unseen indoor environments
II Related Works
II-A Next-Best View and Multisensor Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14]. A common approach uses an active perception formulation to maximize entropy reduction [19, 10]. Ref. [12] combine frontier and sampling based methods to reduce computational overheard for deployment on limited compute UAVs. Ref. [9] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas. Ref. [22] similarly learn an information aware policy offline that proposes waypoints for the robot. These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information. Ref. [46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas. In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings during a search and rescue scenario.
Multisensor exploration involves the inclusion of LiDAR information in planning for visual exploration tasks. State of the art multisensor exploration with multiple UAVs [3] leverages the single agent volumetric exploration algorithm presented in [42]. These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping. Ref. [27] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments. These classifications are utilize by a Pan-Tilt-Zoom camera to perform object detection in Search and Rescue applications. These multisensor approaches demonstrate efficient results in a number of real world environments, however none of these methods is tailored for the indoor environments considered in this article.
II-B World Prediction Based Planning
Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first. Ref. [33] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored. Prediction involves recognizing features common to tunnel networks to improve exploration strategies. Similarly, [36] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers. The planner uses this to explore rooms fully before continuing global exploration. This contrasts from the proposed method that aims to ignore low value regions in order to reduce search completion time. Ref. [31] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context. Ref. [8] learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints. Prediction guide the robot towards likely exit locations.
II-C Lidar Classification
Indoor pointcloud classification is tackled in 3D [39] using the KPConv architecture [37]. This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [38]. Neither work validates on a real robot platform. The proposed work instead focus on low computation 2D scan pixel classification.Ref. [28] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [39], focused on map building rather than contextual labeling for planning. Finally, [34] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping outdoor environments. In contrast LIVES recognizes non-map features for indoor environments.
III Methods
The overall scan classifier and planning module are described in Alg. 1. The key components are detailed in this section.
Algorithm 1 Map-Free LiDAR Informed Search()
$s_{i},x_{i}$
LiDAR Scan, Robot Pose
$x^{*}_{\text{next}}$ (Next viewpoint)
$s_{i}^{\text{class}}=\text{NeuralNetworkClassifier}(s_{i})$
$\mathcal{M}^{\text{LiDAR}}_{i}\leftarrow\text{LiDARMapUpdate}(s_{i},x_{i},\mathcal{M}^{\text{LiDAR}}_{i-1})$
$\mathcal{M}^{\text{Visual}}_{i}\leftarrow\text{VisualMapUpdate}(s_{i},x_{i},\mathcal{M}^{\text{Visual}}_{i-1})$
$\{\mathbb{C}^{\text{LiDAR}}\}=\text{GetLiDARFrontiers}(\mathcal{M}^{\text{LiDAR}}_{i})$
$\{\mathbb{C}^{\text{Visual}}\}=\text{GetVisualFrontiers}(\mathcal{M}^{\text{Visual}}_{i})$
for $n\leftarrow 1$ to $N$ do $\triangleright$ for each frontier $c_{n}$ in $\mathbb{C}$
     for $j\leftarrow 1$ to $4$ do $\triangleright$ for each viewpoint at $c_{n}$
         $\text{util}_{n,j}=\text{ComputeSampleUtility}(c_{n},x_{i},s_{i}^{\text{class}})$
     end for
end for
$x^{*}_{\text{next}}$leftarrow $\arg\max_{n,j}(\text{util}_{n,j})$
return $x^{*}_{\text{next}}$
III-A Environment
The environment is described as a discrete set of points $\mathcal{E}$. This can be split into two subsets $\mathcal{E}^{\text{free}}\subset\mathcal{E}$ and $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$, which represent the free and occupied cells in the environment. Both of these sets are initially unknown to the robot, but are assumed to follow a structure inherent in indoor environments (e.g. hallways, planar walls, open rooms and loops).
III-B Ground Truth LiDAR Scan Classification
The search target of interest belongs to $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$. In particular, scan classification attempts to divide $\mathcal{E}^{\text{occ}}$ into two subsets, namely $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$. In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [7]. LTFs represent the static map obstacles, STFs represent static un-mapped obstacles and DFs represent moving points. Given map information, scan points can be classified into these three categories as follows.
Let $x_{i}$ denote the pose of the robot, and $s_{i}$ denote observation at time step $t_{i}$. Each observation $s_{i}$ consists of $n_{i}$ 2D points, $s_{i}=\{p^{j}_{i}\}_{j=1:n_{i}}$. Observations are transformed from robot local frame into the global frame using an affine transformation $T_{i}\in SE(3)$. Finally, let map $\mathcal{M}$ be represented as a set of lines $\{l_{i}\}_{1:n}$.
III-B1 LTF
First, an analytic ray cast is performed [6] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$. Given observations, the probability that points correspond to one of the lines of that static map can be written
$P(p_{i}^{j}|x_{i},M)=\text{exp}\left(-\frac{\text{dist}(T_{i}p_{i}^{j},l_{j})^{2}}{\Sigma_{s}}\right)$ (1)
where $\Sigma_{s}$ is the scalar variance of observations, which comes from sensor accuracy. If Eq. 1 is greater than a threshold, point $p_{i}^{j}$ is classified as a LTF.
III-B2 STF
Remaining points will be classified as STF or DF. Observations at current time $i$, $p_{i}^{j}$, are compared with prior observations at time $k$, $p_{k}^{l}$ to determine correspondence between points in subsequent observations. The likelihood of the remaining points corresponding to the same point as in a previous laserscan is computed as
$P(p_{i}^{j},p^{l}_{k}|x_{i},x_{k})=\text{exp}\left(-\frac{||T_{i}p_{i}^{j}-T_{k}p^{l}_{k}||^{2}}{\Sigma_{s}}\right)$ (2)
where $p^{l}_{k}$ is the nearest point from $p_{i}^{j}$ among points which does not belong to LTF at other timesteps, defined as
$p^{l}_{k}=\arg\text{min}||T_{i}p_{i}^{j}-T_{k}p^{l}_{k}||$ (3)
When Eq. 2 is greater than some threshold, point $p_{i}^{j}$ is classified as an STF. Remaining points in $p_{i}^{j}$ are classified as DFs. The result is classified scan $s^{\text{class}}_{i}$.
III-C Map-free LiDAR Scan Classification
The classification method defined in III-B is limited by the necessity of static map information $\mathcal{M}$. Therefore, this work uses supervised learning to train a model to reproduce classifications for $\{p^{j}_{i}\}_{j=1:n_{i}}$ in scan $s_{i}$.
III-C1 Dataset and Data Acquisition
The collected dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. $N$ is the total number of data points collected. Collecting a dataset using a real robot requires a human expert to control the robot over long periods of time. This approach is difficult to scale due to mechanical limits of the robot and difficulty of control over large, potentially out-of-sight places. To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot. Human operators easily maneuver this cart to generate the dataset without strain to the robot. We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification. Finally, a laptop is used to maintain the pose estimate [7], collecting the data and generating online ground-truth labels for the dataset. This laptop records 1) robot pose at each timestep, 2) raw scan data, and 3) classified scan from the ground-truth module for dataset $\mathcal{D}$. The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online. Training data can be found on the project website.
III-C2 Architecture and Training
We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification. This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels. During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k$ sliding-window. The history buffer is then concatenated with the updated pixel-wise exponential weighted average. At the initialization of inference, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is initialized. In practice there are a large number of non-map points due to the size of scans ($n_{i}=897$), therefore we apply filtering steps like pooling and removing those in previously inspected areas before these points are used for planning. Code for training a model can be found on the project website.
Figure 2: Pixel-wise LiDAR scan classification model architecture. $k$ is the length of the history buffer. (a) $[3,k]$ pose history matrix containing $[x,y,\theta]$. (b) $[n_{i},k]$ LiDAR range history matrix. (c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average. (d) The model consists of three temporal-convolutional encoders (TCN), one for each input. The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan. In (e) a threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
The full model architecture is shown in Fig. 2. The policy consists of a temporal-convolutional encoder (TCN) for each input. Each encoder contains a single convolutional layer with scan-wise circular padding and a single linear layer. The scan and label encoders have kernel size $[k,k]$, and the pose encoder has kernel size $[1,3]$. Hyperbolic tangent activations are used for all layers. The output of each encoder is the same size as the LiDAR scan to be classified. The output of the pose encoder is summed with the output of the scan encoder as a pose correction, and then normalized by applying a hyperbolic tangent function. The resulting LiDAR scan encoding is multiplied element-wise with the label encoding. A threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point. The policy is trained with mean squared error loss against ground truth classifications. Training for 20 epochs on a workstation Nvidia 3080 12Gb takes 3 minutes.
III-D Map Updates
Two global occupancy maps (search maps) are maintained. The first is a LiDAR global search map and the second is a visual global search map. Global search maps are implemented in this work using ROS Costmap2D [24]. All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$. The robot is equipped with a visual sensor with a cone shaped FoV, implemented as a triangular costmap. The LiDAR sensor has a 360 degree FoV, represented as a square-type costmap. Observations for both sensors are a set of range estimates to the nearest occupied point $e\in\mathcal{E}^{\text{occ}}$. Points between the robot and the nearest occupied point along each ray of the LiDAR scan are free points $e\in\mathcal{E}^{\text{free}}$. Given incoming observations from the robot, each global costmap is updated as free space $e=0$ or occupied $e=1$. This process is shown in Fig. 1(d).
III-E Multisensor Frontier Exploration
The planning module seeks to generate a viewpoint at each planning step that results in the highest likelihood to find the target of interest. This is achieved in several steps: candidate viewpoints are generated, viewpoints are scored, and finally the point with the highest score is fed to the robot for navigation to that point.
III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [44], shown in Fig. 1(e). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Any group of frontier points beyond a minimum size becomes its own frontier. Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering. Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor.
III-E2 Viewpoint Selection
Candidate viewpoints are selected based on a computationally efficient heuristic utility function. High value viewpoints are those that are 1) near the robot, 2) allow inspection of many unknown map points and 3) result in inspection of non-map points. This process is shown in Fig. 1(f). The utility functions is a weighted sum of the following:
  • Penalize distance from robot position to viewpoint
  • Reward expected number of unknown cells discovered
  • Reward viewpoints with frontiers near the path to the candidate viewpoint
  • Reward viewpoints that result in inspection of a scaled number non-map points
Scaling parameters on each term are tuned during experiments. Weight are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered. The next waypoint is selected by sending the highest utility viewpoint, $x^{*}_{\text{next}}$, to the robot.
IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on real robot hardware in unseen indoor environments.
IV-A LiDAR Scan Classification
Figure 3: Ablation studies over policy architecture (top) and injected noise during training (bottom). The inclusion of the label history buffer yields 11.63% higher test accuracy. The policy is robust up to 30% to bit-flipping errors in the label history buffer. The mean accuracy depicts the 5-step moving average.
First we compare the learned classification policy to the ground truth classifier. The output of the policy is a vector of size $n_{i}$ normalized between -1 and 1 for each scan point. If this value is non-negative, the point is classified as a member of $\mathcal{E}^{\text{map}}$. Otherwise the point is classified as a member of $\mathcal{E}^{\text{non-map}}$. We express performance in terms of per-scan accuracy of the classified LiDAR scans from the neural network estimator versus ground truth scan classification from map information. This metric for a classified scan $s^{\text{class}}_{i}$ at time $i$ compared to ground truth $s^{\text{class, true}}_{i}$ is given by
$\text{accuracy}=\frac{1}{{n_{i}}}\sum^{n_{i}}_{j=0}\{+1\text{ if }p^{j}_{i}==p^{j,\text{true}}_{i}\text{ else }0\}$ (4)
The final test accuracy of the policy compared to ground truth is $86.19\%\pm 0.03\%$ with no a priori map requirement. We perform two ablation studies, modulating the model architecture and injected noise during training, with results in Fig. 3. The inclusion of the label encoder is responsible for a 11.63% increase in test accuracy. Because the robot moves relatively slowly compared to the policy update frequency, recently predicted labels strongly bias the current estimate toward the correct classification. We fix the history buffer to 9 time steps, corresponding to 1.8 seconds. The history buffer enables the policy to reason about its environment even though no reference map is provided. During training, uniform random noise is injected into the label history buffer to simulate inference-time auto-regressive accumulated errors. The policy is robust up to 30% of injected noise before the policy begins to over-fit and performance degrades. We select a policy trained with 10% injected noise to deploy on the robot. All ablation studies were performed with a fixed seed (0) and initialized with identical weights. We do not tune the random seed.
IV-B Simulation and Baseline Comparison
Simulations are performed using Gazebo-ROS on the Turtlebot [32]. ROS move base is used to navigate to the goal points given from the viewpoint selection methods compared [25]. Baselines are: 1) a lightly modified version of [42] (Multisensor Frontier Exploration), 2) the Next-Best View Planner (NBVP) [5], 3) a Rapidly-Exploring Random Tree (RRT) [41], and 4) LIVES given ground-truth scan classifications. Ref. [42] described a highly efficient planner for multisensor 3D exploration in underwater environments that has also been extended to multi-drone exploration [3]. The modifications from the original paper are as follows. First, the method is projected into 2D. Second, candidate viewpoint selection is changed. Originally, candidate viewpoints are selected surrounding frontier centroids, but in this work the candidate viewpoints are selected at centroids, with candidate views in four directions. This modified version of the multisensor frontier method is equivalent to the proposed approach if our method is not given any contextual information from the scan classifier.
Map Target NBVP RRT Multisensor Frontier Ground Truth LIVES
Name Difficulty [5] [41] Exploration [42] Classification (Our Method)
Apartment Easy $112(70\%)$ $72(80\%)$ 36 (100%) 34(100%) 34(100%)
Hard $177(30\%)$ $128(60\%)$ 106 (100%) 92(100%) 90(100%)
Office Easy $70(80\%)$ $54\pm(80\%)$ 34 (100%) 26(100%) 24(100%)
Hard $300(0\%)$ 111 (80%) $141(90\%)$ 99(100%) 105(100%)
Table I: Average time (s) for each method to find the object of interest in the given simulation environment.
Two Gazebo world environments are tested with easy and hard setups in each. A figure depicting them is shown on the project website. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES. The easy setting incorporates fewer objects in the environment while hard has more objects in the environment. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position. The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods. In the easy apartment setting we observe similar performance among the three planners. However in the hard setting, where LIVES and Ground-Truth perform nearly identically, there is roughly a 15% improvement with the addition of the knowledge of non-map points compared with the Multisensor method. There exists a large gap in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant. In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%. Overall, the results indicate that the proposed method to inform the NBV planner with non-map information outperforms the baselines across the full spectrum of environments. We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
Figure 4: Ground truth maps of the two experimental environments. (a) shows the 15mx38m hallway environment and (b) shows the 20mx30m apartment.
Figure 5: A time lapse of a simple experiment scenario. In (a) the robot pose is a centered on the visual costmap. The global search map, frontier points and non-map points are also shown. The target object is behind the robot starting position. The result of the first planning step is also shown. The planner takes advantage of the concentration of non-map points behind the robot. Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).
IV-C Hardware Experiments
Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig. 4. The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. A Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera is utilized. The search target of interest is one or two small rolling suitcases and YOLOv5 [18] is used for detection of the object. Objects are localized by first converting PointCloud data into the optical frame and then by conversion to pixel coordinates using camera intrinsics. A time lapse of one experiment in the apartment is shown in Fig. 5 where the search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side. However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby. Finally, we compare the accuracy of the trained classifier in these out of distribution environments to ground truth information. The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
V Conclusions
We present a method for search in unseen environments using contextual information to inform planning. Results show LIVES to outperform several baseline methods in varied environments by 10-30%. Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme. The proposed approach is deployed in an unseen environment to certify the learning method and visual planner for real-world scenarios. Ongoing work aims to extend the LIVES planner to multiple robots.